### v1

# thanks: Author contributions: J.E. and T. A. contributed equally to this workthanks: Author contributions: J.E. and T. A. contributed equally to this work

Abstract
We consider the effect of insoluble surfactant on jet drop formation by bubble bursting. We demonstrate quantitative agreement on the jet velocity and drop size between laboratory experiments and numerical simulations by using the measured surface tension dependence on surfactant concentration as the equation of state for the simulations. We demonstrate that surfactants increase the drop size for bubbles without precursor capillary waves, which is opposite in trend to bubbles with precursor capillary waves, due to the effect of Marangoni stresses on the focusing of the collapsing cavity.
When bubbles burst at a liquid-air interface, the resulting cavity collapses and the associated capillary waves focus to emit a liquid jet. The jet can then pinch to form drops, referred to as jet drops, which contribute to ocean aerosol emissions [1, 2], health [3, 4], and industrial aroma [5]. Many applications involve complex solutions such as viscoelastic fluids [6], contaminated interfaces with oil [7, 8, 9], or surface-active agents (surfactants) [10, 11, 12, 13].
Jet drop production has been studied extensively [14, 15]. The bursting of a bubble of radius $R_{0}$ for a Newtonian liquid of density $\rho$, dynamic viscosity $\mu$, with constant surface tension $\gamma_{c}$, and gravitational constant $g$ can be described by two non-dimensional numbers (assuming that the external air flow is negligible with high liquid-to-gas density and viscosity ratios),
$\textit{La}=\frac{\rho\gamma_{c}R_{0}}{\mu^{2}},~{}\textit{Bo}=\frac{\rho gR_{0}^{2}}{\gamma_{c}},$ (1)
where the Laplace number La denotes a balance between capillary and viscous forces, and the Bond number Bo denotes a balance between capillary and gravitational forces. The Laplace number is related to the Ohnesorge number Oh via $\textit{La}=\textit{Oh}^{-2}$ and can be considered a Reynolds number. It is known that for $\textit{La}\gtrsim\textit{La}_{*}\approx 500$, jet drops are ejected [16, 17, 18, 19, 20]. The jetting dynamics are controlled mainly by La, with Bo controlling the initial bubble shape and modifying the jetting dynamics for $\textit{Bo}\geq 0.1$ [21, 19, 22, 23, 24, 20]. For $\textit{La}>\textit{O}(10^{4})$, the capillary waves due to the collapsing cavity contain many precursor waves (ripples), whereas for $\textit{La}=\textit{O}(10^{3})$, the precursor waves are damped by viscosity and the collapsing cavity can focus more efficiently, producing self-similar dynamics that give a faster and narrower jet [25, 26, 5, 27], leading to smaller jet drops.
In practice, interfaces are often contaminated with surfactants from biological or anthropogenic origin. The surfactant surface concentration $\Gamma$ affects the surface tension of the interface via some equation of state $\gamma=\gamma(\Gamma)$ (often referred to as the surface tension isotherm) [28]. The resulting surface tension gradient gives rise to a stress: the Marangoni stress. Jet drop production depends significantly on the surface tension and therefore is affected by surfactants, as shown experimentally and numerically [10, 11, 29, 13, 12]. Previous studies [10, 11, 13, 12] investigated surfactant effects for $\textit{La}\in[2\cdot 10^{4},10^{5}]$, where precursor capillary waves are present, and found that adding surfactants decrease the ejected drop radii. Indeed, at high La, surfactants allow the cavity to focus more since the Marangoni stress damps the precursor waves, leading to a faster and narrower jet.
Figure 1: Surfactant effects on jet drops. (a) Laplace number $La_{d}=\rho\gamma_{c}R_{d}/\mu^{2}$ of the first ejected drop radius $R_{d}$ as a function of La. Experimental points (filled) are colored by the bulk Triton X-100 concentration. Numerical points (open) are colored by the surfactant parameter $\beta$ (Eq. (3)). The black line and shaded area represent the average and typical spread found in the data (without surfactant) [20, 18]. Triangles represent experimental [11, 13] (filled) and numerical [10, 12] (open) data available in the literature on jet drops with surfactant. Arrows indicate the effect of adding surfactant at a given La. (b) Experimentally measured surfactant equation of state (markers) for several bulk surfactant concentrations as a function of the surface surfactant concentration $\Gamma$ (same colors as (a), 50% glycerol), with the corresponding fits (dashed curves) according to Eq. (3). (c,d) Illustrations of the jet drop dynamics and comparisons between experiments and simulations ((c,d): $(\textit{La},\textit{Bo},\Delta\gamma_{\infty},E)=(2000,0.16,0.55,0.5)$, (c): $\beta=0.24$, (d): $\beta=0.37$). The background images are experimental snapshots and the color [30] overlays show the interface profile and the surfactant concentration from the simulations. Supplementary videos are available.
Quantitative comparisons between experiments and numerical simulations have remained limited [31] as experiments report the surfactant bulk concentration without having access to spatial and temporal variations in surface concentration and the surfactant equation of state is not always available. Separately, two-phase Navier-Stokes simulations with surfactant transport and Marangoni stresses enable detailed access to the surfactant dynamics. However, connecting the multiple numerical parameters influencing Marangoni stresses to practical surfactant effects has remained challenging.
In this Letter, we show, through experiments and simulations, that for bubbles with efficient self-similar cavity collapse that is not controlled by precursor capillary waves ($\textit{La}=\textit{O}(10^{3})$), the ejected drop radii increase upon the addition of insoluble surfactants (Fig. 1(a)) by changing the shape of the self-similar collapsing cavity. We present a systematic approach to quantitatively relate experiments and simulations. We use experimentally measured surface tension isotherms $\gamma=\gamma(\Gamma)$ as the equation of state in the simulation (Fig. 1(b)), and obtain excellent agreement between the experimental and simulation results in terms of the temporal and spatial evolution of the cavity collapse, jet formation, and drop ejection (Fig. 1(c, d)).
The structure of the Letter is as follows: we first describe, and then quantitatively compare the numerical and experimental frameworks. Then, the effects of surfactants on jet drop production at low La are described. Finally, we propose a physical mechanism for the trend reversal, dependent on La based on how Marangoni stresses modulate the cavity collapse.
Our numerical simulations solve the 2D axisymmetric two-phase Navier-Stokes equations using the coupled level-set/volume-of-fluid method [32, 33] using the Basilisk open-source library [34, 35]. Air-water density and viscosity ratios are used. The numerical configuration is adapted from Berny et al. [20, 36], who considered bubble bursting without surfactants. The Marangoni stress is resolved with an integral formulation for the surface tension [33] and the insoluble surfactant concentration is transported via a coupled phase-field/volume-of-fluid method [37, 38] (see supplementary material).
Just after bubble bursting, the liquid-air interface is initialised using the Young-Laplace equations [39, 40, 41] to compute the theoretical shape of the cavity, with the bubble cap removed at the foot of the cap $r=r_{\text{cap}}$ (see Fig. 1(c)) [26, 5, 20, 10, 12]. We consider a uniform initial surfactant concentration $\Gamma=\Gamma_{0}$, and test the sensitivity to surfactant accumulation on the cap film, which retracts into the tip at $r=r_{\text{cap}}$ upon bursting, represented by an initial non-uniform concentration
$\Gamma(s_{\text{cap}})=\Gamma_{0}\left(1+ER_{0}\delta(s_{\text{cap}})\right),$ (2)
[…]
Our experiments are performed by generating bubbles in a small tank and recording their bursting at the surface; see previous experiments with similar setups [5, 11, 44]. The bubble lifetimes vary from $0-16$ s. The bubbles are generated using a syringe pump pushing air through a small capillary, resulting in sizes $R_{0}\approx 1.0$ mm. We use three different solutions (see table 1), one consisting only of deionized water, one with 33% by weight of glycerol and one with 50% glycerol. For each of the solutions, various amounts of Triton X-100, a nonionic surfactant, are added, in concentrations ranging from 1 to 20% of the critical micelle concentration (CMC). This surfactant adsorbs or desorbs from an interface slowly: the typical timescale ranges from minutes to hours [45, 46, 47, 48], allowing us to assume insolubility on capillary timescales ($O(1~{}{\rm ms})$).
wt % La Bo $\Delta\gamma_{\infty}$ $\beta$
50 2 000 - 2 700 0.15 - 0.17 0.55 0.1 - 0.3
33 8 800 - 12 000 0.12 - 0.16 0.58 0.1 - 0.4
0 67 000 - 83 000 0.12 - 0.14 0.52 0.0 - 0.4
Table 1: The non-dimensional parameters in the experiment, for the three glycerol weight fractions used [49, 50]. The uncertainty of La is mostly due to the temperature (estimated as $21\pm 1^{\circ}$), which affects the dynamic viscosity $\mu$. See supplementary material for the dimensional parameters.
To quantitatively compare experiments and simulations in addition to the trends, we measure the surface tension isotherm (equation of state) experimentally, which is substituted into the simulations. In a Langmuir trough, the surface tension is measured using the Wilhelmy plate method, where two barriers compress the surfactant-covered surface, reducing the area $A$ available for surfactant molecules. The surfactant is considered insoluble during the measurement (again, the adsorption and desorption are slow [48]), and a change in surface area therefore corresponds to a change in surface concentration $\Gamma$ through $A/A_{0}=\Gamma_{0}/\Gamma$, where $A_{0}$ and $\Gamma_{0}$ are the area and surface concentration in the initial state, respectively. Curves obtained through this procedure are shown (markers) in Fig. 1(b). The solutions that we used may have minor external contamination since the 0% CMC case with 50% glycerol is not flat in Fig. 1(b).
We fit the isotherms with a model that takes into account the surfactant physics while minimizing the number of parameters (dashed curves in Fig. 1(b)). In particular, we model the surface tension saturation that is experimentally visible at large $\Gamma$. Such saturation is important for the jet drop dynamics since the focusing of the capillary waves during cavity collapse leads to a much larger surfactant concentration at jet formation than the initial concentration (see Fig. 1(c,d)). For simplicity, we use a two-parameter $\tanh$ model [31, 37]:
$\gamma/\gamma_{c}=1-\Delta\gamma_{\infty}\tanh\left(\frac{\beta}{\Delta\gamma_{\infty}}\frac{\Gamma}{\Gamma_{0}}\right)$ (3)
with two nondimensional parameters $\Delta\gamma_{\infty}$ and $\beta$.
Physically, $\Delta\gamma_{\infty}$ is the nondimensional difference in surface tension between a clean and saturated interface since Eq. 3 gives $\gamma/\gamma_{c}=1$ at $\Gamma=0$ and $\gamma/\gamma_{c}\approx 1-\Delta\gamma_{\infty}$ for $\Gamma\gg\Gamma_{0}$. Therefore, $\Delta\gamma_{\infty}$ is constant for a given solution. In this Letter, only the experimentally-derived $\Delta\gamma_{\infty}$ is used, but the simulations show that there is sensitivity to $\Delta\gamma_{\infty}$ (see supplementary material). The surfactant parameter $\beta$ can be interpreted as the slope of the isotherm since $\gamma/\gamma_{c}\approx 1-\beta(\Gamma/\Gamma_{0})$ for $\Gamma\ll\Gamma_{0}$.
The simulation is sensitive to the particular choice of the function used to model the two-parameter isotherm (Eq. 3). While this ambiguity can be avoided by fitting the isotherm more closely using more parameters, which gives a comparable performance in describing the dataset, the use of more parameters leads to interpretability and overfitting issues (discussed in more details in the supplementary material).
We investigate the effect of surfactants on the first ejected jet drop for low La, systematically varying $\beta$ and comparing with the experimental case with 50% glycerol. Sensitivity to La and non-uniformity in the initial surfactant distribution is considered so as to compare with the experimental variations in viscosity and bubble lifetime.
We focus on three dependent variables $(R_{d},V_{d},t_{d})$ accessible both experimentally and numerically, where the first ejected drop radius is $R_{d}$, $V_{d}$ is the velocity of the first ejected drop (at pinch off), and $t_{d}$ is the time between bubble rupture and jet pinch off as a function of the surfactant parameter $\beta$. In Fig. 2 the experimental points (solid) are colored by the bubble lifetime prior to bubble rupture. The shape of the numerical points (open) are given by $E$ and are colored by La. Fig. 2(a,b,c) show that as $\beta$ increases, the drop radius $R_{d}$ increases, drop velocity $V_{d}$ decreases, and pinch-off time $t_{d}$ increases for both experiments and simulations. Increasing $E$ or decreasing $La$ in the simulations results in $R_{d}$ and $t_{d}$ increasing and $V_{d}$ decreasing. The experiment shows a spread for each measured $\beta$. Such a spread is physical and arises from the variation in bubble lifetime, whose effect seems to be similar to increasing the excess surfactant $E$, which suggests that variations in accumulation of surfactant at the top of the bubble, as captured by the parameter $E$, may be an explanation for lifetime effects [51, 43, 52]. Indeed, when considering relationships implicit in $\beta$ (see Fig. 2(d,e,f)), the experimental and numerical points collapse onto a curve with strong agreement between the experimental and numerical data. Such agreement suggests that the jet dynamics are being resolved accurately by the simulation. Physically, the collapse onto a curve suggests that the jet dynamics may be independent of the surfactant; it is the cavity collapse that is modulated by the surfactant, as Marangoni stresses affect regions of high curvature where surfactant accumulates during the cavity collapse. Again, it should be noted that the exact degree of agreement between the simulation and experiment is sensitive to the particular model of the isotherm used.
Figure 2: Effect of the surfactant parameter $\beta$ on the first ejected drop radius $R_{d}$, velocity $V_{d}$, and pinch-off time $t_{d}$. (a,b,c) Relationships between $(R_{d},V_{d},t_{d})$ and $\beta$, nondimensionalised by the bubble radius $R_{0}$, capillary velocity $\gamma_{c}/\mu$, and inertiocapillary timescale $\sqrt{\rho R_{0}^{3}/\gamma_{c}}$ respectively. (d,e,f) Relationships between non-dimensionalised $(R_{d},V_{d},t_{d})$. The same numerical and experimental data points are used throughout (a-f). For $\textit{Bo}=0.16$ as considered, $r_{\text{cap}}/R_{0}\approx 0.5$ and hence results with $E=0-1$ are shown.
We show that the size amplification due to the surfactants occurs since Marangoni stresses disturb the self-similar collapse of the cavity shape, leading to a slower and thicker jet formation. Indeed, without surfactants, at low $\textit{La}=1000-5000$ [27], it has been shown that the interface profile is self-similar just before and after jet production [25, 26, 5, 27] at time $t_{0}$ with a $|t-t_{0}|^{\frac{2}{3}}$ scaling.
Fig. 3 shows the cavity collapse ($t<t_{0}$) in physical (a) and self-similar (b) coordinates, and the jet production ($t>t_{0}$) in physical (c) and self-similar (d) coordinates. A typical clean bubble (gray) is compared to a corresponding contaminated bubble (colored) at low $\textit{La}=2400$. The clean case in Fig. 3 is the analogue to Figs. 3,4 of Lai et al. [27] ([27] did not consider surfactants).
Figure 3: Self-similar profile comparison between a clean ($\textit{La}=2400,\textit{Bo}=0.16$) and correspondingly contaminated ($\textit{La}=2400,\textit{Bo}=0.16,\beta=0.3,\Delta\gamma_{\infty}=0.55,E=0.5$) bubble. The clean case is analogous to Figs. 3,4 of [27]. The interface is given by $z=h(r,t)$. (a) The time evolution of the cavity collapse in time $(t<t_{0})$ for the clean (gray) and contaminated bubble (colors). The times shown are $(t_{0}-t)/t_{ic}=0.018,0.024,0.03$, where $t_{ic}=\sqrt{\frac{\rho R_{0}^{3}}{\gamma_{c}}}$ is the inertia-capillary timescale. (b) The time evolution of the cavity collapse in self-similar coordinates, corresponding to curves in (a). (c) The time evolution of jet production ($t>t_{0}$) (same convention as (a)), at times $(t-t_{0})/t_{ic}=0.008,0.011,0.015,0.019$. (d) The time evolution of jet production in self-similar coordinates corresponding to curves in (c). In (b,d), $\min$ is taken over the profile of the cavity or jet. The colors for (a,b,c,d) are given by the surfactant concentration $\Gamma$ and Marangoni stress $d\gamma/ds$ for surface tension $\gamma$ and arclength $s$ measured from $r=0$. The particular choice of times plotted in (a,b,c,d) is the same as [27], which captures the self-similarity.
As noted in the literature, for clean bubbles, at low La, the cavity forms a sharp corner prior to jet formation (see Fig. 3(a,b)), which results in high capillary pressure that leads to a fast and thin jet (see Fig. 3(c,d)). In comparison, as can be seen in Fig. 3(a,b), the case with surfactants has a smoother corner, which results in a smaller magnitude capillary pressure and leads to a slower and thicker jet (see Fig. 3(c,d)). Physically, there is an accumulation of surfactants (see Fig. 3(a,b)) near the corner of the cavity (such accumulation has also been seen in [10, 12] for high La), which results in a low surface tension region leading to Marangoni stress that smooths the corner. Mathematically, it has been shown that the possible self-similar interface profile [53] for inertio-capillary cavity collapse is a family parametrized by the far-field (with respect to the local region of cavity collapse) cone angle and interfacial flow field. Thus, the cavity shape can be interpreted as the result of Marangoni flow that acts to select a smoother corner.
Separately, for $\textit{La}\gg 5000$, multiple capillary waves appear as precursor waves ahead of the dominant wave (referred to as the corner above). In such a regime, the details of the precursor waves play a key role in the jet behavior, and hence the drop amplification mechanism identified in this Letter, is not applicable. Indeed, previous studies have suggested that for $\textit{La}\gg 5000$, the Marangoni stress damps the capillary waves in such a way that the drop radius decreases [10, 11, 12, 13]; which we also observe in our numerical and experimental data (see Fig. 1(a) and supplementary material).
In summary, we demonstrate that the trend reversal of drop radius with La (when adding surfactant), first an increase and then decrease, is due to the change in the dominant physical mechanism that produces the jet drop at low La (high capillary pressure of the cavity corner) and high La (precursor wave). We have identified that the Marangoni stress acts to increase the ejected drop radius for low La, distinct from the high La regime, where the ejected drop radius decreases instead. These arguments are supported by the experimental observations (see Fig. 1(a)), which show that the transition occurs near $\textit{La}\approx 10^{4}$.
As a direct extension, it would be interesting to include how other physical effects, such as surfactant solubility and viscoelasticity, would affect the trend reversal. Considering a larger class of non-uniform initial surfactant distributions may also be important. Finally, many Marangoni flow problems would likely benefit from the approach taken in this Letter, where experimentally-derived surface tension isotherms are used in the numerical simulations.
Acknowledgements.
This work was supported by NSF grant 2242512 to L.D. This work used TAMU Faster at Texas A&M High Performance Research Computing through allocation OCE140023 to L.D. from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program [54], which is supported by U.S. National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296.